OpenStudy (anonymous):

What is the equation of the circle with center (–4, –3) that passes through the point (6, 2)?

OpenStudy (anonymous):

Help please?

OpenStudy (anonymous):

the general form for a circle of center \((h,k)\) and radius \(r\) is \[(x-h)^2+(y-k)^2=r^2\] you have \((h,k)=(-4,-3)\) so the first part should be ok

OpenStudy (anonymous):

what you need is \(r^2\) which is the square of the distance between \((-4,-3)\) and \((6,2)\) use the distance formula without the square root to find it

OpenStudy (anonymous):

you know how to find the distance between two points?

OpenStudy (anonymous):

nevermid i got the answer wrong

OpenStudy (anonymous):

wrong answers are good too`